OpenStudy (dylanx):

Find the real x-values: (x)^3 - 3x-1 = 0

OpenStudy (anonymous):

Sum and product of roots of a cubic I think. \[ax^3 + bx^2 + cx + d=0\] \[p+q+r=\frac{-b}{a}\] \[p+q+r=0\] \[pqr=\frac{-d}{a}\] \[pqr=\frac{1}{1}\]

OpenStudy (anonymous):

p, q and r are the x-values. \[pq+pr+qr = \frac{c}{a}\] \[pq+pr+qr = -3\]

OpenStudy (anonymous):

I think you need to use trial and error instead.

OpenStudy (anonymous):

Refer to the Mathematica attachment.